OpenStudy (anonymous):

Solve log11 (y + 8) + log11 4 = log11 60.

OpenStudy (anonymous):

@phi

OpenStudy (mathstudent55):

First, use this on the left side: \(\log a + \log b = \log (ab)\)

OpenStudy (mathstudent55):

Then use this: If \(\log a = \log b\) then \(a = b\).

OpenStudy (anonymous):

y=28 check right @mathstudent55

OpenStudy (anonymous):

@mathstudent55 how do you solve this equation for answers between 0-2Pie \[3\tan ^{4}\alpha= 1+\sec ^{2}\alpha \]

OpenStudy (anonymous):

y=28 check right

OpenStudy (anonymous):

@Michele_Laino is it right check

OpenStudy (welshfella):

for the trig equation i suggest you use the identity sec^2 a = 1 + tan^2 a

OpenStudy (anonymous):

y=28 am i right check

OpenStudy (anonymous):

yes you are right with y=28

OpenStudy (welshfella):

log 4(y + 8) = log 60 4y + 32 = 60 y = ?

OpenStudy (welshfella):

not 28

OpenStudy (anonymous):

4y+32=60 4y+32-32=60-32 4y=28 y=28 i'm right

OpenStudy (anonymous):

@Michele_Laino i'm right see

OpenStudy (welshfella):

4y = 28 not y

OpenStudy (anonymous):

you forgot to divide by 4

OpenStudy (welshfella):

y = 28/4

OpenStudy (anonymous):

no fraction

OpenStudy (michele_laino):

please what is 28/4=...? @sallyfield888

OpenStudy (mathstudent55):

\(\log_{11} (y + 8) + \log_{11} 4 = \log_{11} 60\) \(\log_{11} (y + 8)4 = \log_{11} 60\) \(\log_{11} (4y + 32) = \log_{11} 60\) \(4y + 32 = 60\) \(4y = 28\) \(y = 7\)

OpenStudy (anonymous):

7

OpenStudy (mathstudent55):

@sallyfield888 Correct!

OpenStudy (michele_laino):

that's right! @sallyfield888

OpenStudy (anonymous):

7 is right

OpenStudy (welshfella):

3 tan^4 a = 1 + 1 + tan^2 a 3tan^4 a - tan^2 a - 2 = 0 you need to solve this equation

OpenStudy (welshfella):

if you let t = tan^2 a we get 3t^2 - t - 2 = 0 (3t + 2)(t - 1) = 0 so find t t = tan^ a

OpenStudy (welshfella):

from the values of tan ^ 2 a you can then find a

OpenStudy (welshfella):

tan^2 a cannot be negative for real a so you cane ignore the negative root

OpenStudy (anonymous):

im confused on how you went from the equatoin to the t =tan^2a

OpenStudy (mathstudent55):

For this equation, let's look at a trig identity you need \(3\tan ^{4}\alpha= 1+\sec ^{2}\alpha\) Start here: \(\sin^2 \alpha + \cos^2 \alpha = 1\) \(\dfrac{\sin^2 \alpha}{\color{red}{\cos^2 \alpha}} + \dfrac{\cos^2 \alpha}{\color{red}{\cos^2 \alpha}} = \dfrac{1}{\color{red}{\cos^2 \alpha}}\) \(\tan^2 \alpha + 1 = \sec^2 \alpha\) Now use this substitution in your equation: \(3\tan ^{4}\alpha= 1+\color{red}{\sec ^{2} \alpha}\) \(3\tan ^{4}\alpha= 1+\color{red}{\tan^2 \alpha + 1}\) \(3\tan ^{4}\alpha =\tan^2 \alpha + 2\) \(3\tan ^{4}\alpha - \tan^2 \alpha - 2 = 0\) Now you have a quadratic in \(tan^2 \alpha \) You can solve it directly for \(\tan^2 \alpha\), or you can make a substitution. Let's solve it for \(\tan^2 \alpha\) without making a substitution. \((3\tan ^2\alpha + 2 )(\tan^2 \alpha - 1) = 0\) \(3\tan ^2\alpha + 2 = 0\) or \(\tan^2 \alpha - 1 = 0\) \(3\tan ^2\alpha = -2\) or \(\tan^2 \alpha = 1\) Left equation gives no solution since a^2 = negative has no real solutions. \(\tan^2 \alpha = 1\) \(\tan \alpha = \pm \sqrt 1 \) \(\tan \alpha = \pm 1\) In degrees: \(\alpha = 45^o \) or \(\alpha = 135^o\) or \(\alpha = 225^o \) or \(\alpha = 315^o\) In radians: \(\alpha = \dfrac{\pi}{4} \) or \(\alpha = \dfrac{3 \pi}{4}\) or \(\alpha = \dfrac{5 \pi}{4}\) or \(\alpha = \dfrac{7 \pi}{4}\)

OpenStudy (anonymous):

thank you so much, that helped alot!!